# Statistical Inference for Temporal Difference Learning with Linear Function Approximation

### v1

@@ Appendix D Experimental details
D.1 The Markov decision process and the feature map
@@
From the expressions above, we remark that, the values of $q$ and $V^{\star}(s)$ with $s\geq d$ are fixed for all $\bm{q}\in\mathcal{Q}$ which is ensured by the construction (112). It is shown in Theorem 2 of Li et al., 2024b that the sequence $\{\mathcal{M}_{\bm{q}}\}_{q\in\mathcal{Q}}$ is hard to distinguish for a considerably large set $\mathcal{Q}$.
D.2 Parameter specification
@@
    In order to illustrate the Berry-Esseen bound as indicated by Theorems 3.2 and 3.3, we measure the difference between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$ by
@@
    In order to illustrate the efficiency of our variance esitmator $\widehat{\bm{\Lambda}}_{T}$, we measure the averaged Frobenius norm of the difference between $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ among $100$ independent trials. As is indicated by A.12, this is a direct proxy for the TV distance between the zero-mean Gaussian distributions with the two corresponding variance matrices.
@@
    In order to validate our statistical inference procedure, we run Algorithm 1 on the MDP with $d=3$ using $\alpha=\frac{2}{3}$, construct both individual and simultaneous confidence intervals for $\bm{\theta}^{\star}$, and measure the coverage rates of these confidence intervals in $1000$ independent trials. Results are shown in section 4.2.
D.3 Full experimental results
D.3.1 High-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$
@@
$O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{d}{\delta}}\right)$
@@
D.3.2 Berry-Esseen bound
Figure 6 illustrates how the finite-sample distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ converges to its asymptotic Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$. Notably, plot (a), which is a replication of Figure 2(b) in Section 4, indicates that the convergence is slower when $\alpha=\frac{1}{2}$. This directly contradicts the suggestion by Samsonov et al., (2024) to choose $\alpha=\frac{1}{2}$ in order to yield a faster in-distribution convergence of the TD estimation error. In fact, with $d=3$ fixed and other choices of $\alpha$, the convergence rate are similar when $T$ is large. Meanwhile, plot (b) shows the convergence generally is slower when $d$ is larger.
@@ D.3.2 Berry-Esseen bound
D.3.3 Estimator of the covariance matrix
@@
Figure 7: The decay of the difference bewteen estimated and asymptotic covariance matrices, measured in squared Frobenius norm, when (a) $d=3$, for different choices of $\alpha$; and (b) $\alpha=\frac{2}{3}$, for different choices of $d$.

### v3

@@ Appendix D Experimental details
D.1 Online algorithm to construct the variance estimator
1:MDP transition kernel $\mathcal{P}$, stationary distribution $\mu$, discount factor $\gamma$, feature map $\bm{\phi}$, initial stepsize $\eta_{0}$, stepsize decay parameter $\alpha\in(0.5,1)$, sample size $T$
2:Initialize $\bm{\theta}=\bm{0}$, $\overline{\bm{\theta}}=\bm{0}$, $\overline{\bm{A}}=\bm{0}$, $\overline{\bm{AA}}=\bm{0}$, $\overline{\bm{Ab}}=\bm{0}$, $\overline{\bm{bb}}=\bm{0}$
3:for $t=0:(T-1)$ do
4:     Sample $s\sim\mu$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s)$
5:     Compute:
$\bm{A}\leftarrow\bm{\phi}(s)[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})],\quad\bm{b}\leftarrow r(s)\bm{\phi}(s)$
6:     Perform TD update:
$\bm{\theta}\leftarrow\bm{\theta}-\eta_{0}t^{-\alpha}(\bm{A}\bm{\theta}-\bm{b})$
7:     Perform online updates:
$\overline{\bm{\theta}}\leftarrow\overline{\bm{\theta}}+\frac{1}{t}(\bm{\theta}-\overline{\bm{\theta}})$
$\overline{\bm{A}}\leftarrow\overline{\bm{A}}+\frac{1}{t}(\bm{A}-\overline{\bm{A}})$
$\overline{\bm{AA}}\leftarrow\overline{\bm{AA}}+\frac{1}{t}(\bm{A}\otimes\bm{A}-\overline{\bm{AA}})$
$\overline{\bm{Ab}}\leftarrow\overline{\bm{Ab}}+\frac{1}{t}(\bm{A}\otimes\bm{b}+\bm{b}\otimes\bm{A}-\overline{\bm{Ab}})$
$\overline{\bm{bb}}\leftarrow\overline{\bm{bb}}+\frac{1}{t}(\bm{b}\otimes\bm{b}-\overline{\bm{bb}})$
8:end for
9:Construct $\widehat{\bm{\Gamma}}$:
$\widehat{\bm{\Gamma}}\leftarrow\mathsf{reshape}(\overline{\bm{AA}}(\overline{\bm{\theta}}\otimes\overline{\bm{\theta}})-2\overline{\bm{Ab}}(\overline{\bm{\theta}})+\overline{\bm{bb}},(d\times d))$
10:Construct the variance estimator:
11:Averaged TD estimator $\overline{\bm{\theta}}$, covariance matrix estimator $\widehat{\bm{\Lambda}}$
The basic properties of Kronecker products indicate
Since Algorithm 1 iterates from $t=0$ to $t=T-1$ in a single loop and does not store the samples for each iteration, it has a space complexity is $O(Td^{4})$ and a space complexity of $O(d^{4})$. Here, the $d^{4}$ term stems from calculating the Kronecker product $\bm{A}\otimes\bm{A}$.
D.2 The Markov Decision Process and the feature map
@@
From the expressions above, we remark that, the values of $q$ and $V^{\star}(s)$ with $s\geq d$ are fixed for all $\bm{q}\in\mathcal{Q}$ which is ensured by the construction (117). It is shown in Theorem 2 of Li et al. (2024) that the sequence $\{\mathcal{M}_{\bm{q}}\}_{q\in\mathcal{Q}}$ is hard to distinguish for a considerably large set $\mathcal{Q}$.
D.3 Parameter specification
@@
    In order to illustrate the Berry-Esseen bound as indicated by Theorems B.3 and B.4, we measure the difference between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$ by
@@
    In order to illustrate the efficiency of our variance estimator $\widehat{\bm{\Lambda}}_{T}$, we measure the averaged Frobenius norm of the difference between $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ among $100$ independent trials. As is indicated by A.14, this is a direct proxy for the TV distance between the zero-mean Gaussian distributions with the two corresponding variance matrices.
@@
    In order to validate our statistical inference procedure, we run Algorithm 1 on the MDP with $d=3$ using $\alpha=\frac{2}{3}$, construct both individual and simultaneous confidence intervals for $\bm{\theta}^{\star}$, and measure the coverage rates of these confidence intervals in $1000$ independent trials. Details to follow in section D.4.4.
D.4 Full experimental results
D.4.1 High-probability convergence guarantee for the TD estimator
@@
$O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}\right)$
@@
D.4.2 Berry-Esseen bound
Figure 4 illustrates how the finite-sample distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ converges to its asymptotic Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$. Notably, plot (a), which is a replication of Figure 2 in Section 4, indicates that the convergence is slower when $\alpha=\frac{1}{2}$. This directly contradicts the suggestion by Samsonov et al. (2024) to choose $\alpha=\frac{1}{2}$ in order to yield a faster in-distribution convergence of the TD estimation error. In fact, with $d=3$ fixed and other choices of $\alpha$, the convergence rate are similar when $T$ is large. Meanwhile, plot (b) shows the convergence generally is slower when $d$ is larger.
@@ D.4.2 Berry-Esseen bound
D.4.3 Estimator of the covariance matrix
@@
Figure 5: The decay of the difference between estimated and asymptotic covariance matrices, measured in squared Frobenius norm, when (a) $d=3$, for different choices of $\alpha$; and (b) $\alpha=\frac{2}{3}$, for different choices of $d$.
D.4.4 Confidence intervals
Finally, we validate our statistical inference framework on the MDP with $d=3$ and TD learning with $\alpha=\frac{2}{3}$. We run 1000 independent trials of Algorithm 1, and at each iteration $T$, we construct both the individual and the simultaneous confidence intervals. Specifically, the individual confidence intervals for the three entries of $\bm{\theta}^{\star}$ are constructed by